OpenStudy (anonymous):

Limit as x approach 0+ of √x lnx

OpenStudy (anonymous):

can you use l'hopital?

OpenStudy (anonymous):

is it \[\sqrt{x}\times \ln (x)\] or \[\sqrt{x \times \ln x}\]

OpenStudy (anonymous):

usual gimmick it to rewrite as \[\frac{\ln(x)}{\frac{1}{\sqrt{x}}}\] and then use l'hopital

OpenStudy (anonymous):

\[\sqrt{x} lnx \]

OpenStudy (anonymous):

I've tried with L'hospital but am getting stuck in the work

OpenStudy (anonymous):

When I use L'hopital I get \[\lim_{x \rightarrow 0+} = -2\times \sqrt{x} = 0\] and when I graph it on my TI I get the same

OpenStudy (anonymous):

thank you :)